OpenStudy (meowlover17):

Which statement is true about the sum square root of 50 + square root of 6? A,B,C,D Answer Choices

OpenStudy (meowlover17):

It is rational and equal to 2. It is rational and equal to 8. It is irrational and equal to 2square root of 2. It is irrational and equal to 8square root of 2.

OpenStudy (anonymous):

\[\sqrt{50}+\sqrt{6}=\sqrt{25*2}+\sqrt{6}=5\sqrt{2}+\sqrt{6}\]

OpenStudy (meowlover17):

Sorry i think i asked the wrong question...

OpenStudy (meowlover17):

\[\sqrt{50}+\sqrt{18}\]

OpenStudy (meowlover17):

What statement is true about the sum above...

OpenStudy (anonymous):

\[=5\sqrt{2}+\sqrt{9*2}=5\sqrt{2}+3\sqrt{2}=\sqrt{2}(5+3)=8\sqrt{2}\]

OpenStudy (anonymous):

It;s D

OpenStudy (meowlover17):

Thanks

OpenStudy (meowlover17):

Medal earned

OpenStudy (anonymous):

YES ! Another question?

OpenStudy (meowlover17):

Thanks for your Help M_lowreen one more

OpenStudy (anonymous):

post it